OpenStudy (anonymous):

Need to put into scientific notation: 7.0 * 10 negpower4 + 6.2 * 10power2.

OpenStudy (anonymous):

if the answer is 620.0007, then how is that written in sn?

OpenStudy (anonymous):

6.200007 * 10^2

OpenStudy (anonymous):

So is that your final answer?

OpenStudy (anonymous):

yes, Regis :)

OpenStudy (anonymous):

thxs;)

OpenStudy (loser66):

I don't understand, please explain me, @Euler271

OpenStudy (anonymous):

hmmm

OpenStudy (loser66):

@satellite73 hmmmm???

OpenStudy (anonymous):

scientific notation or the Regis Philbin joke?

OpenStudy (anonymous):

\[7\times 10^{-4}\times 6.2\times 10^2\]?

OpenStudy (loser66):

@Euler271 surely the problem.

OpenStudy (anonymous):

its just another way of writing it. 6.2 * 10^2 = 6.2 * 100 = 620

OpenStudy (anonymous):

\[7\times 6.2=43.4\] and \[10^{-4}\times 10^2=10^{-2}\] so i get \[43.4\times 10^{-2}=.434\]

OpenStudy (loser66):

@satellite73 It's not multiplication, it's addition

OpenStudy (anonymous):

and any number can be written this way. number.evertyhingelse * 10^exponent

OpenStudy (anonymous):

ooh addition doh then i guess you have to add

OpenStudy (anonymous):

true

OpenStudy (loser66):

\[7.0^10^{-4} = 0.0007\\6.2*10^2= 620\\620+ 0.0007= 620.0007 \]why we have *10^2 after them?

OpenStudy (anonymous):

it was 6.20007 * 10^2 it's just another way of writing it. onenumber.restofnumbers * 10^exp

OpenStudy (loser66):

don't we have to put it in scientific notation and don't we have to follow significant figures?

OpenStudy (anonymous):

the equation is 7.0 x 10^-4 + 6.2 x 10^2

OpenStudy (anonymous):

does'nt specify to keep proper sig figs. if that's the case then we'd ignore then 7*10^-4 part

OpenStudy (anonymous):

yes, specifically I need to know how to solve as sn

OpenStudy (anonymous):

so, your final answer still stands, Meredith? dith?

OpenStudy (anonymous):

I want to put this to rest, so is 6.20007 x 10^2 correct?